# Michael Schuckers Professor of Data Science and Sports Analytics School of Data Science, UNC Charlotte

### v1

@@ Abstract
Using data on 103 recent P4 college football hires, we built a statistical model for predicting a coach’s success at their new school. For each hire, we collected data about their background and experiences, the previous success as a head coach or coordinator and their success since hiring. Over 50 variables on these factors were recorded though we used 29 of these in building our predictive model. Our measure of success is based upon Bill Connelly’s SP+ team ratings relative to the performance on the same metric of the school in the 15 year prior to their selection as head coach. Using a cross-validated regularized linear regression, we obtain a predictive model for coaching success. Among the important factors for predicting a successful hire are having been a previous college head coach, having won a prior conference championship as a head coach, leaving a job as an Offensive Coordinator, age and quality of the hiring school’s team in the previous 15 years. While we do find these factors are important for the prediction of a successful coaching hire, the trends here are weak. With 66% accuracy, the model does identify coaching hires that will outperform team performance in the 15 years before the hire. However, no combination of these factors leads to high predictability of identifying a successful coaching hire. All of the data and code for this paper are available in a Github repository.
@@ 1 Introduction
College football is big business. The changes due to the use of player’s name, image and license (NIL) that allow players to be openly (and appropriately) paid for their services, has brought a new influx of funds. A December 2024 study by CNBC using fiscal year 2023 numbers estimated that the top university athletic programs are valued at over $1$ Billion dollars and have annual revenue of over $200$ Million, [cnbc_college_worth]. Unsurprisingly the executive that leads the college football operation, the head coach, are well compensated. In a 2023 study, David Evans found that ”80% of the Highest-Earning Public Employees are College Head Coaches.” [highest_state]
@@ 3 Data Summaries
Table 1: Top Coaching Hires by Sp+ Rel.
Name Hiring School SP+ Rel
Josh Heupel UCF 22.6
Rhett Lashlee SMU 17.5
Sonny Dykes SMU 16.1
Mike Gundy Oklahoma State 13.8
Mike Elko Duke 10.7
Matt Campbell Iowa State 9.9
Dabo Swinney Clemson 9.7
Bret Bielema Illinois 9.4
Lane Kiffin Ole Miss 8.7
Josh Heupel Tennessee 8.5
@@ 4 Model and Results
$\widehat{SP+Rel}=-0.042$ $+0.463\mbox{ Was College HC before at any level}$
@@
The largest factors in our model were whether or not the coach’s previous job was as an OC (PreviousOC), whether they had ever been a college coach previously (prevCollegeHC) and the teams performance over the last 15 years (team sp prev). Having been hired from a OC position, is associated with about a $3.7$ point increase in a coach’s SP+ Rel. While if a hire was previously a college football HC, the predicted increase in their team’s performance increases by nearly half a point. Age is a factor in our model that is negatively correlated with coaching success though the effect is very small, about a hundredth of a point per game for an additional 10 years of age. This seems quite negligible. For a newly hired 50 year old coach on a teams who’s SP+ averaged 0 in the last 15 years with none of the other characteristics in the model, we expect their SP+ Rel to be -0.042 or just slightly below what the team had done in the previous 15 years. Just below The last term in our prediction equation involves a team’s average SP+ Rel over the previous 15 years and that is negatively associated with how a coaching hire’s teams will perform. We can interpret the value here of 0.306-0.306 as the expected decrease in a team’s average performance for every additional point of SP+ Rel that the hiring team had over the previous 15 years. This is a regressive effect, since SP+ for an average team should be around zero, then for a new coaching hire good teams we would expect them to get worse and very good teams we would expect them to get much worse, while for below average teams we would expect them to get better just by making the hire. One possible explanation for this is a regression to the mean as team performance tends to move to the average of all teams, which for SP+ is around zero. Another explanation is college football has moved to an era of more parity and relative coaching performance is not immune to moving with that trend.
@@
The hires with the ten highest and ten lowest predicted SP+ Rel are listed in Table 2. We can see that the model suggests that the hire with the highest predicted success was Rhett Lashlee at SMU. Eight of the ten top predicted SP+ Rel values had positive observed values. Tony Elliott at Virginia and Jedd Fisch at Arizona were the exceptions. Elliott has improved his Virginia teams and, if he stays, may yet get to positive territory. Fisch left Arizona after a successful second season. Among the coaches predicted to have negative SP+ Rel, nine of the ten are in or have finished in negative territory. The exception is Kirby Smart who has impressively elevated Georgia to new heights. Looking at the bottom ten, most of the schools there are considered to be ‘elite’ or nearly so. The model predicts that succeeding at these schools is harder and
@@
Rhett Lashlee SMU 17.5 6.2
Josh Heupel UCF 22.6 5.6
@@
Mike Gundy Oklahoma State 13.8 4.1
@@
Tony Elliott Virginia -6.4 3.5
Mike Locksley Maryland 0.5 2.9
@@
Lance Leipold Kansas 8.4 2.3
Willie Taggart Oregon -9.1 -5.6
Mark Richt Miami -2.4 -5.7
Luke Fickell Wisconsin -14.8 -5.8
@@
Kirby Smart Georgia 4.2 -6.4
@@
Brent Venables Oklahoma -8.5 -6.8
Ed Orgeron LSU -4.8 -7.3
@@ 5 Discussion
Much has been made about the recent success of Curt Cignetti and he has done an amazing job as the HC at Indiana. So it is worth mentioning that Cignetti is not in the data we used for modelling because we have less than two full years of him coaching at Indiana. Regardless, Cignetti is an outlier. It is possible to regenerate all of the figures in this analysis with Cignetti and other coaches hired after the 2023 season and it is very clear which observation corresponds to the exceptional job Cignetti has done at Indiana. That said, an outlier is not a trend and the signal from these data is weak. Hiring a coach is hard. There are no magic attributes, at least not among the factors we investigated, that guarantee future success. The data as can be seen in Figures 1 through 5 is very noisy and the trends, while tempting, are mild. Great coaching are hard to find and some schools will find themselves waiting for their next great coach  [^3] Just like Vladimir and Estragon.
@@ 6 Appendix
Table 3: Bottom Coaching Hires by Sp+ Rel.
@@
Luke Fickell Wisconsin -14.8

### v2

# Waiting for Dabo: [^1] With apologies to Samuel Beckett. A machine learning model for predicting Power 4 college football coaching hire success

@@ Abstract
Using data on 103 recent P4 college football hires, we built a statistical model for predicting a coach’s success at their new school. For each hire, we collected data about their background and experiences, the previous success as a head coach or coordinator and their success since hiring. Over 50 variables on these factors were recorded though we used 29 of these in building our predictive model. Our measure of success is based upon Bill Connelly’s SP+ team ratings relative to the performance on the same metric of the school in the 15 year prior to their selection as head coach. Using a cross-validated regularized linear regression, we obtain a predictive model for coaching success. Among the important factors for predicting a successful hire are having been a previous college head coach, leaving a job as an Offensive Coordinator, age and quality of the hiring school’s team in the previous 15 years. While we do find these factors are important for the prediction of a successful coaching hire, the trends here are weak. With 66% accuracy, the model does identify coaching hires that will outperform team performance in the 15 years before the hire. However, no combination of these factors leads to high predictability of identifying a successful coaching hire. All of the data and code for this paper are available in a Github repository.
@@ 1 Introduction
College football is big business. The changes due to the use of player’s name, image and license (NIL) that allow players to be openly (and appropriately) paid for their services, has brought a new influx of funds. A December 2024 study by CNBC using fiscal year 2023 numbers estimated that the top university athletic programs are valued at over $1$ billion US dollars and have annual revenue of over $200$ million, [cnbc_college_worth]. Unsurprisingly the executive that leads the college football operation, the head coach, are well compensated. In a 2023 study, David Evans found that “80% of the Highest-Earning Public Employees are College Head Coaches.” [highest_state]
@@ 2 Data
Table 1: Best Performing Coaching Hires by Sp+ Rel.
Name Hiring School SP+ Rel
Josh Heupel UCF 22.6
Rhett Lashlee SMU 17.5
Sonny Dykes SMU 16.1
Mike Gundy Oklahoma State 13.8
Mike Elko Duke 10.7
Matt Campbell Iowa State 9.9
Dabo Swinney Clemson 9.7
Bret Bielema Illinois 9.5
Lane Kiffin Ole Miss 8.8
Josh Heupel Tennessee 8.7
@@ 4 Model and Results
$\widehat{SP+Rel}=-0.054$ $+0.493\mbox{ Was College HC before at any level}$
$+3.750\mbox{ Previous job was as an OC}$
$-0.002\mbox{ (Age - 50)}$
@@
The largest factors in our model were whether or not the coach’s previous job was as an OC (PreviousOC), whether they had ever been a college coach previously (prevCollegeHC) and the teams performance over the last 15 years (team sp prev). Having been hired from a OC position, is associated with about a $3.75$ point increase in a coach’s SP+ Rel. While if a hire was previously a college football HC, the predicted increase in their team’s performance increases by nearly half a point. Age is a factor in our model that is negatively correlated with coaching success though the effect is very small, about two-hundredth of a point per game for an additional 10 years of age. This seems quite negligible. For a newly hired 50 year old coach on a teams who’s SP+ averaged 0 in the last 15 years with none of the other characteristics in the model, we expect their SP+ Rel to be -0.054 or just slightly below what the team had done in the previous 15 years. The last term in our prediction equation involves a team’s average SP+ Rel over the previous 15 years and that is negatively associated with how a coaching hire’s teams will perform. We can interpret the value here of 0.307-0.307 as the expected decrease in a team’s average performance for every additional point of SP+ Rel that the hiring team had over the previous 15 years. This is a regressive effect, since SP+ for an average team should be around zero, then for a new coaching hire good teams we would expect them to get worse and very good teams we would expect them to get much worse, while for below average teams we would expect them to get better just by making the hire. One possible explanation for this is a regression to the mean as team performance tends to move to the average of all teams, which for SP+ is around zero. Another explanation is college football has moved to an era of more parity and relative coaching performance is not immune to moving with that trend.
@@
The hires with the ten highest and ten lowest predicted SP+ Rel are listed in Table 2. We can see that the model suggests that the hire with the highest predicted success was Rhett Lashlee at SMU. Eight of the ten top predicted SP+ Rel values had positive observed values. Tony Elliott at Virginia and Jedd Fisch at Arizona were the exceptions. Elliott has improved his Virginia teams and, if he stays, may yet get to positive territory. Fisch left Arizona after a successful second season. Among the coaches predicted to have negative SP+ Rel, eight of the ten are in or have finished in negative territory. The exceptions are Dan Mullen whose team performance dropped significantly over his four years at Florida and Kirby Smart who has impressively elevated Georgia to new heights. Six of these ten coaches have already been fired with only Mullen, Smart, Fickell and Venables still employed at these schools. Venables’ Oklahoma 2025 team is having their best season with him as a HC and so there is hope for a turnaround there. Looking at the bottom ten schools, most of the schools there are considered to be ‘elite’ or nearly so. The model predicts that succeeding at these schools is hard partly because they are already at the upper tail of the distribution. Overall the model accurately predicts whether a coach will have a positive or negative SP+ Rel 67% of the time. Table 5
@@
Rhett Lashlee SMU 17.8 6.3
Josh Heupel UCF 22.6 5.7
@@
Mike Gundy Oklahoma State 13.8 4.2
@@
Tony Elliott Virginia -5.6 3.5
Mike Locksley Maryland 0.4 2.9
@@
Lance Leipold Kansas 8.5 2.3
Dan Mullen Florida 1.1 -5.7
Mark Richt Miami -2.4 -5.8
Luke Fickell Wisconsin -15.0 -5.9
@@
Kirby Smart Georgia 4.5 -6.4
@@
Brent Venables Oklahoma -8.9 -6.8
Ed Orgeron LSU -4.8 -7.4
@@ 5 Discussion
Much has been made about the recent success of Curt Cignetti and he has done an amazing job as the HC at Indiana. So it is worth mentioning that Cignetti is not in the data we used for modelling because we have less than two full years of him coaching at Indiana. Regardless, Cignetti is an outlier. It is possible to regenerate all of the figures in this analysis with Cignetti and other coaches hired after the 2023 season and it is very clear which observation corresponds to the exceptional job Cignetti has done at Indiana. That said, an outlier is not a trend and the signal from these data is weak. Hiring a coach is hard. There are no magic attributes, at least not among the factors we investigated, that guarantee future success. The data as can be seen in Figures 1 through 5 is very noisy and the trends, while tempting, are mild. Great coaching are hard to find and many schools find themselves waiting for their next great coach.
@@ 6 Appendix
Table 3: Worst Performing Coaching Hires by Sp+ Rel.
@@
Luke Fickell Wisconsin -15.0
@@
Table 5: Truth Table for SP+ Rel
actual
SP+ Rel<0<0 SP+ Rel>0>0
predicted $\widehat{SP+Rel}<0$ 51 23
$\widehat{SP+Rel}>0$ 11 18